# Prompt Optimization via Retrieved Reasoning Assets and Multi-Agent Analysis

### v1

@@
Abstract.
Prompt optimization has emerged as an effective alternative to retraining for improving the performance of Large Language Models (LLMs). However, most existing approaches treat evaluation as a black box, relying solely on numerical scores while offering limited insight into why a prompt succeeds or fails. They also depend heavily on trial-and-error refinements, which are difficult to interpret and control. In this paper, we introduce MA-SAPO, a Multi-Agent framework for Score-Aware Prompt Optimization. Compared to prior methods, MA-SAPO explicitly couples evaluation outcomes with structured reasoning to guide systematic edits. The framework specifically consists of two stages: during the Reasoning Phase, agents collaboratively explain metric scores, diagnose weaknesses, and synthesize targeted refinements that are stored as reusable reasoning assets; during the Test Phase, agents retrieve these assets to analyze optimized prompts and apply only evidence-grounded edits. By turning evaluation signals into interpretable reasoning chains, MA-SAPO produces prompt refinements that are more transparent, auditable, and controllable. Experiments on the HelpSteer1/2 benchmarks demonstrate consistent improvements over single-pass prompting, retrieval-augmented baselines, and prior multi-agent strategies, validating the effectiveness of our approach. [^1] Code is available at: https://anonymous.4open.science/r/MA-SAPO-F313.
Automatic Prompt Optimization, Score-Aware Prompt Optimization, Retrieval-Augmented Generation, Large Language Models (LLMs), Multi-Agent System
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Multi-agent systemsccs: Information systems Query representationccs: Computing methodologies Natural language generationccs: Information systems Relevance assessment
Figure 1. Comparative overview of prompt optimization methods. MA-SAPO enhances interpretability by linking evaluation scores to reasoning-driven and evidence-based refinements.
1. Introduction
@@
Large Language Models (LLMs) have emerged as powerful tools capable of tackling a wide range of tasks, from reasoning and summarization to complex dialogue and code generation (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023). However, their performance remains highly sensitive to the wording, structure, and examples embedded within prompts (Pryzant et al., 2023; He et al., 2024; Errica et al., 2024; Cao et al., 2024; Razavi et al., 2025). Consequently, prompt optimization has rapidly emerged as a practical alternative for improving model behavior, as it avoids the need for costly retraining or parameter updates.
@@
Early research explored single-pass prompt optimization strategies that treated the model itself as the optimizer. For instance, methods such as Chain-of-Thought (CoT) prompting (Wei et al., 2022), role assignment through carefully designed system instructions (Wu et al., 2023; Kong et al., 2023), or structured variants such as Tree-of-Thought (ToT) (Yao et al., 2023), Graph-of-Thought (GoT) (Besta et al., 2024), and step-back prompting (Zheng et al., 2023) demonstrated that reasoning patterns or role signals embedded into prompts could substantially improve downstream performance. While effective, these approaches typically operate in isolation: a single model instance refines prompts without incorporating systematic feedback from prior attempts or external resources.
@@
As LLMs have advanced, multi-agent frameworks have gained traction as a way to diversify perspectives and coordinate multiple roles in the optimization process (Zhou et al., 2025; Zhang et al., 2025; Han et al., 2025). By assigning diverse specialized agents to different tasks such as decomposition, evaluation, or refinement, these systems move beyond single-pass rewriting to more structured optimization pipelines. At the same time, optimization-theoretic methods (Li et al., 2025b; Cui et al., 2025) and label-efficient pipelines (Xiang et al., 2025; Dong et al., 2025) have sought to reduce the reliance on costly feedback while ensuring scalability across tasks.
Despite these advances, several common limitations remain: (1) most existing frameworks reduce evaluation to outcome scores, treating it as a black box and leaving practitioners uncertain about why certain prompts succeed or fail; (2) optimization often relies on repetitive trial-and-error refinements that consume significant computation while offering little transparency; (3) even when reasoning is incorporated, it usually remains implicit and is not transformed into explicit, auditable artifacts that can guide systematic edits; (4) current pipelines largely optimize for higher outcome scores but offer limited interpretability and controllability, preventing users from understanding trade-offs or applying targeted adjustments.
@@
To address these limitations, we propose MA-SAPO, a sequential reasoning framework for prompt optimization that explicitly links metric outcomes to actionable edits through structured reasoning artifacts. Specifically, MA-SAPO operates in two complementary phases. In the reasoning phase, three agents transform annotated scores into progressively enriched artifacts: (1) a Metric Explainer Agent interprets evaluation dimensions, (2) a Diagnostician Agent uncovers error sources and trade-offs, and (3) an Action Synthesizer Agent produces concrete edit directives. These artifacts are stored as reusable reasoning assets. In the test phase, a retrieval-augmented pipeline leverages these assets: (1) an Analyzer Agent contrasts the current prompt with retrieved exemplars to identify improvement opportunities, while (2) a Refiner Agent applies targeted edits supported by diagnostic evidence, ensuring that optimizations are interpretable and controllable.
@@
Extensive experiments on the open-source HelpSteer1/2 benchmarks (Wang et al., 2023, 2024a) demonstrate that MA-SAPO consistently outperforms single-pass prompting, retrieval-augmented generation, and prior multi-agent baselines, highlighting the importance of structured reasoning as a bridge between evaluation and optimization. Our main contributions include:
@@
  • We propose a score-aware multi-agent training pipeline that distills evaluation outcomes into reusable, semi-structured reasoning assets, enabling transparent prompt explanations, fine-grained diagnoses, and actionable edit directives.
@@
  • We design MA-SAPO, a modular and computationally efficient framework that achieves interpretable, auditable, and controllable improvements across diverse datasets and model backbones, while substantially reducing token and API call budgets.
@@
  • We conduct extensive experiments on HelpSteer1 and HelpSteer2 benchmarks, demonstrating consistent gains across multiple evaluation metrics and outperforming six state-of-the-art baselines in both effectiveness and efficiency.
2. Related Work
@@
Single-Pass Prompt Optimization. Prompt optimization has rapidly evolved as a practical means of improving the performance of LLMs without retraining (Wei et al., 2022; Pryzant et al., 2023; Sun et al., 2023; Li et al., 2025a). Early studies highlighted the centrality of prompts as control variables for steering model behavior. Specifically, few-shot prompting (Brown et al., 2020) demonstrated that models could adapt to unseen tasks using in-context examples alone, while retrieval-augmented generation (RAG) (Lewis et al., 2020; Lee et al., 2025) injected external knowledge to improve factuality and adaptability. These two paradigms, context construction and evidence injection, laid the foundation for single-pass prompt optimization methods. Subsequent research explored reasoning-augmented prompting. Chain-of-Thought (CoT) (Wei et al., 2022) introduced intermediate reasoning steps. Building on this idea, Tree-of-Thought (ToT) (Yao et al., 2023) generalized linear reasoning chains into branching structures with self-evaluation via lookahead and backtracking. Other structured strategies, such as Skeleton-of-Thought (SoT) (Ning et al., 2023), separate planning from execution to parallelize generation, while Graph-of-Thought (GoT) (Besta et al., 2024) organizes reasoning trajectories into graph topologies. Evolutionary and heuristic single-pass approaches have also been explored. For instance, PromptBreeder (Fernando et al., 2023) evolves a population of prompts through mutation and selection, while PromptWizard (Agarwal et al., 2024) refines prompts via feedback-driven critique-and-synthesis loops.
@@
Multi-Agent Based Prompt Optimization. Beyond single-pass methods, multi-agent systems have emerged as a powerful paradigm for prompt optimization (Du et al., 2023; Liu et al., 2024; Chen et al., 2024; Li et al., 2024). Early agentic frameworks such as Self-Ask (Press et al., 2022) introduced decomposition by generating sub-questions and integrating external retrieval, laying the groundwork for cooperative prompting strategies. Also, Multi-Agent Debate (MAD) (Du et al., 2023) advanced this idea by allowing agents to iteratively present and critique answers in a roundtable format, thereby improving factual grounding and reasoning quality. Reflection-based frameworks (Shinn et al., 2023; Khan et al., 2024) further enhanced adaptability by storing feedback as memory for subsequent optimization attempts. Building on this foundation, recent systems formalized agent roles and workflows to optimize prompts more systematically (Li et al., 2023). Subsequent work explicitly targeted prompt optimization as the objective. For instance, MASS (Zhou et al., 2025) searches over agent topologies while performing block-level prompt refinement, followed by global optimization. Additionally, MARS (Zhang et al., 2025) leverages a Planner and Teacher–Critic–Student dialogue to iteratively propose, critique, and revise prompts in structured rounds, and also MAPGD (Han et al., 2025) integrates pseudo-gradient feedback with bandit-style exploration, offering sample-efficient optimization through agent collaboration. However, most existing frameworks treat evaluation as a black box with limited interpretability and rely on costly trial-and-error refinements. MA-SAPO addresses this by grounding optimization in score-aware reasoning assets that link evaluation outcomes to actionable, interpretable, and auditable refinements.
Figure 2. Overview of the MA-SAPO framework. The training phase constructs reasoning assets $(\mathcal{R}_{i}=(\mathcal{C}_{i},\mathcal{D}_{i},\mathcal{E}_{i}))$ from annotated prompt–response pairs, while the test phase retrieves top-k training examples and their reasoning assets to refine a new test prompt.
3. Methodology
3.1. Preliminaries
We denote the training dataset as
@@ 3.1. Preliminaries
where only prompts are available, and corresponding responses are generated during evaluation. Our objective is to generate structured reasoning assets from $\mathcal{D}_{train}$ that can later be retrieved and utilized to optimize new prompts and their responses. Thus, the training phase serves as data generation, while the test phase operates as a retrieval-augmented optimization pipeline.
@@
3.2. Training Phase: Reasoning Asset Construction
@@
3.2.1. Metric Explainer Agent $G_{\text{exp}}$
The Metric Explainer Agent provides natural language justifications for the given scores. Formally, given $\tau_{i}$, the agent produces a reasoning card $\mathcal{C}_{i}$:
(1) $\mathcal{C}_{i}=G_{\text{exp}}(\tau_{i},P_{\text{exp}}),$
@@ 3.2.1. Metric Explainer Agent $G_{\text{exp}}$
3.2.2. Diagnostician Agent $G_{\text{diag}}$
@@
(2) $\mathcal{D}_{i}=G_{\text{diag}}(\tau_{i},\mathcal{C}_{i},P_{\text{diag}}).$
@@
3.2.3. Action Synthesizer Agent $G_{\text{syn}}$
@@
(3) $\mathcal{E}_{i}=G_{\text{syn}}(\tau_{i},\mathcal{C}_{i},\mathcal{D}_{i},P_{\text{syn}}),$
@@
As a result, $\mathcal{R}_{i}$ is stored as the reasoning assets aligned with $\tau_{i}$, forming the retrieval corpus for the test phase. Importantly, these reasoning assets are designed as semi-structured text, which makes them machine-parseable and directly usable by downstream agents in the test phase.
3.3. Test Phase: Retrieval and Reformulation
@@
3.3.1. Analyzer Agent $G_{\text{ana}}$
@@
(4) $\mathcal{A}=G_{\text{ana}}(p_{\text{test}},\{\tau_{i},\mathcal{R}_{j}\}_{j=1}^{k},P_{ana}).$
@@
3.3.2. Refiner Agent $G_{\text{ref}}$
@@
(5) $\hat{p}=G_{\text{ref}}(p_{\text{test}},\mathcal{A},P_{\text{ref}}).$

### v2

# MA-SAPO: Multi-Agent Reasoning for Score-Aware Prompt Optimization

@@
Abstract
Prompt optimization has become a practical way to improve the performance of Large Language Models (LLMs) without retraining. However, most existing frameworks treat evaluation as a black box, relying solely on outcome scores without explaining why prompts succeed or fail. Moreover, they involve repetitive trial-and-error refinements that remain implicit, offering limited interpretability or actionable guidance for systematic improvement. In this paper, we propose MA-SAPO: a new Multi-Agent Reasoning for Score Aware Prompt Optimization framework that links evaluation outcomes directly to targeted refinements. Specifically, in the Training Phase, multiple agents interpret evaluation scores, diagnose weaknesses, and generate concrete revision directives, which are stored as reusable reasoning assets. In the Test Phase, an analyzer agent retrieves relevant exemplars and assets for a new prompt, and a refiner agent applies evidence-based edits to improve the prompt and its response. By grounding optimization in structured reasoning, MA-SAPO ensures edits are interpretable, auditable, and controllable. Experiments on the HelpSteer1/2 benchmarks show that our framework consistently outperforms single-pass prompting, retrieval-augmented generation, and prior multi-agent methods across multiple evaluation metrics.
1 Introduction
@@
Large Language Models (LLMs) have emerged as powerful tools capable of tackling a wide range of tasks, from reasoning and summarization to complex dialogue and code generation (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023). However, their performance remains highly sensitive to the wording and examples embedded within prompts (Pryzant et al., 2023; He et al., 2024; Cao et al., 2024; Razavi et al., 2025). Consequently, prompt optimization has rapidly emerged as a practical alternative to costly retraining or parameter updates for improving model behavior.
@@
Early research explored single-pass prompt optimization that effectively treats the LLM as its own optimizer. Approaches such as Chain-of-Thought (CoT) prompting (Wei et al., 2022), role conditioning via carefully crafted system instructions (Wu et al., 2023; Kong et al., 2024), and structured reasoning schemes including Tree-of-Thought (ToT) (Yao et al., 2023), Graph-of-Thought (GoT) (Besta et al., 2024), and step-back prompting (Zheng et al., 2024) demonstrated that encoding explicit reasoning structure or role signals in the prompt can materially improve downstream performance. Nevertheless, these methods are largely self-contained, relying on a single model instance to revise prompts without leveraging systematic feedback across iterations or external resources.
@@
As LLMs have advanced, multi-agent frameworks have emerged to broaden viewpoints and coordinate specialized roles in prompt optimization (Zhou et al., 2025; Zhang et al., 2025; Han et al., 2025). By delegating subtasks such as decomposition, evaluation, and refinement to distinct agents, these systems extend beyond single-pass rewriting into more structured pipelines. In parallel, optimization-theoretic approaches (Li et al., 2025a; Cui et al., 2025) and label-efficient schemes (Xiang et al., 2025; Dong et al., 2025) aim to curb reliance on expensive feedback while scaling across tasks.
Despite these advances, several common limitations remain: (i) evaluation is often collapsed to scalar scores, obscuring why prompts succeed or fail; (ii) refinement remains largely trial-and-error, incurring high computational cost with limited transparency; (iii) reasoning, when used, is typically implicit rather than distilled into auditable artifacts that drive targeted edits; and (iv) pipelines prioritize score gains over interpretability and controllability, which hinders practitioners from analyzing trade-offs and making precise adjustments in real-world applications.
@@
To address these limitations, we introduce MA-SAPO, a sequential reasoning framework for prompt optimization that explicitly maps metric outcomes to actionable edits via structured reasoning artifacts. MA-SAPO has two phases. In the training phase, three agents convert annotated scores into reusable assets: (1) a Metric Explainer interprets evaluation dimensions, (2) a Diagnostician identifies error sources and trade-offs, and (3) an Action Synthesizer produces concrete edit directives. In the test phase, a retrieval-augmented pipeline applies these assets: an Analyzer contrasts the current prompt with retrieved exemplars to surface improvement opportunities, and a Refiner performs targeted edits grounded in diagnostic evidence, yielding interpretable and controllable optimization.
@@
Experiments on HelpSteer1/2 (Wang et al., 2024c; b) show that MA-SAPO consistently surpasses single-pass prompting, retrieval-augmented generation, and prior multi-agent baselines, underscoring structured reasoning as an effective bridge between evaluation and optimization. Our contributions are:
@@
  • A score-aware training pipeline that distills evaluation outcomes into reusable, semi-structured reasoning assets for explanation, diagnosis, and edit directives.
@@
  • MA-SAPO, a modular and compute-efficient framework that delivers interpretable, auditable, and controllable gains while reducing token and API-calls.
@@
  • Extensive results on HelpSteer1/2 demonstrating consistent multi-metric improvements and outperforming six baselines in both effectiveness and efficiency.
2 Related work
2.1 Single-pass prompt optimization
@@
Prompt optimization has become an effective way to improve LLM performance without retraining (Wei et al., 2022; Pryzant et al., 2023; Sun et al., 2023; Li et al., 2025b). Early work established prompts as key control variables for steering model behavior. In particular, few-shot prompting (Brown et al., 2020) showed that models can adapt to new tasks through in-context examples, while retrieval-augmented generation (RAG) (Lewis et al., 2020; Lee et al., 2025) improves factuality by injecting external knowledge. Together, these approaches laid the foundation for single-pass prompt optimization. Later work introduced reasoning-augmented prompting. Chain-of-Thought (CoT) (Wei et al., 2022) added intermediate reasoning steps, and Tree-of-Thought (ToT) (Yao et al., 2023) extended this into branching reasoning with self-evaluation. Other structured variants include Skeleton-of-Thought (SoT) (Ning et al., 2023), which separates planning from execution, and Graph-of-Thought (GoT) (Besta et al., 2024), which models reasoning as graph-structured trajectories. Evolutionary and heuristic methods have also been explored: PromptBreeder (Fernando et al., 2023) evolves prompts through mutation and selection, while PromptWizard (Agarwal et al., 2024) refines them via critique-and-synthesis loops.
2.2 Multi-agent based prompt optimization
@@
Beyond single-pass methods, multi-agent systems have become an important paradigm for prompt optimization (Du et al., 2024; Liu et al., 2024; Chen et al., 2024; Li et al., 2024; Seo et al., 2026). Early agentic approaches such as Self-Ask (Press et al., 2023) introduced question decomposition and retrieval, while Multi-Agent Debate (MAD) (Du et al., 2024) improved reasoning through iterative answer exchange and critique. Reflection-based methods (Shinn et al., 2023; Khan et al., 2024) further enhanced adaptability by incorporating feedback from prior attempts. Building on these ideas, later frameworks structured agent roles and workflows more explicitly (Li et al., 2023) and began targeting prompt optimization itself. For example, MASS (Zhou et al., 2025) searches over agent topologies with local and global prompt refinement, MARS (Zhang et al., 2025) uses Planner and Teacher–Critic–Student interactions for iterative revision, and MAPGD (Han et al., 2025) combines pseudo-gradient feedback with bandit-based exploration. However, many existing methods still treat evaluation as a black box and rely on costly trial-and-error refinement. MA-SAPO instead grounds optimization in score-aware reasoning assets, making revisions more interpretable, auditable, and actionable.
3 Methodology
Figure 1: MA-SAPO pipeline overview. Three training agents convert annotated prompt–response pairs into reusable Reasoning Assets $\mathcal{R}_{i}=(\mathcal{C}_{i},\mathcal{D}_{i},\mathcal{E}_{i})$. At test time, retrieved assets guide the Analyzer $G_{\mathrm{ana}}$ and Refiner $G_{\mathrm{ref}}$ to produce an optimized prompt $\hat{p}$ with higher scores across all five evaluation dimensions.
3.1 Preliminaries
Let the training dataset be denoted as
@@ 3.1 Preliminaries
where only prompts are available, and corresponding responses are generated during evaluation. Our objective is to generate structured reasoning assets from $\mathcal{D}_{train}$ that can be retrieved to optimize new prompts and their corresponding responses. Thus, the training phase serves as data generation, while the test phase operates as a retrieval-augmented optimization pipeline.
@@
3.2 Training phase: reasoning asset construction
@@
3.2.1 Metric explainer agent $G_{\text{exp}}$
The Metric Explainer Agent generates natural language justifications for the assigned scores. Formally, given $\tau_{i}$, the agent produces a reasoning card $\mathcal{C}_{i}$:
$\mathcal{C}_{i}=G_{\text{exp}}(\tau_{i},P_{\text{exp}}),$ (1)
@@ 3.2.1 Metric explainer agent $G_{\text{exp}}$
3.2.2 Diagnostician agent $G_{\text{diag}}$
@@
$\mathcal{D}_{i}=G_{\text{diag}}(\tau_{i},\mathcal{C}_{i},P_{\text{diag}}).$ (2)
@@
3.2.3 Action synthesizer agent $G_{\text{syn}}$
@@
$\mathcal{E}_{i}=G_{\text{syn}}(\tau_{i},\mathcal{C}_{i},\mathcal{D}_{i},P_{\text{syn}}),$ (3)
@@
As a result, $\mathcal{R}_{i}$ is stored as reasoning assets aligned with $\tau_{i}$, forming the retrieval corpus for the test phase. Importantly, these reasoning assets are designed as semi-structured text, which makes them machine-parseable and directly usable by downstream agents in the test phase. For instance, edit directives often follow a pattern such as “add a constraint to …”, which the refiner agent can reliably identify and execute.
3.3 Test phase: retrieval and reformulation
@@
3.3.1 Analyzer agent $G_{\text{ana}}$
@@
$\mathcal{A}=G_{\text{ana}}(p_{\text{test}},\{\tau_{i},\mathcal{R}_{j}\}_{j=1}^{k},P_{\text{ana}}).$ (4)
@@
3.3.2 Refiner agent $G_{\text{ref}}$
@@
$\hat{p}=G_{\text{ref}}(p_{\text{test}},\mathcal{A},P_{\text{ref}}).$ (5)